OpenStudy (anonymous):

Write an equation in standard form of the line that passes through the given point and has the given slope m. (–4, 3); m = 1/2

OpenStudy (mathmate):

Use y=m(x-x0) + y0

OpenStudy (hunus):

\[y-y_{1}=m(x-x_{1})\] \[y_{1}=3\] \[x_{1}=-4\] \[y-3=1/2(x+4)\] \[-x/2+y=5\]